OpenStudy (anonymous):

integrate from 0 to pi/4 dx/1+cos^2(x)

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4} dx/1+\cos^2x\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integral+of+1%2F(1%2Bcos%5E2(x)) Click show steps, I worked this like half way down and then I said no thanks haha